# On the Effectiveness of Large Language Models in Domain-Specific Code Generation

### v3

@@ Abstract.
Large language models (LLMs) such as ChatGPT have shown remarkable capabilities in code generation. Despite the great achievement, they rely on enormous training data to acquire a broad spectrum of open-domain knowledge. Besides, their evaluation revolves around open-domain benchmarks like HumanEval, which primarily consist of programming contests. Therefore, it is hard to fully characterize the intricacies and challenges associated with particular domains (e.g., web, game, and math). In this paper, we conduct an in-depth study of the LLMs in domain-specific code generation. Our results demonstrate that LLMs exhibit sub-optimal performance in generating domain-specific code, due to their limited proficiency in utilizing domain-specific libraries. We further observe that incorporating API knowledge as prompts can empower LLMs to generate more professional code. Based on these findings, we further investigate how to efficiently incorporate API knowledge into the code generation process. We experiment with three strategies for incorporating domain knowledge, namely, external knowledge inquirer, chain-of-thought prompting, and chain-of-thought fine-tuning. We refer to these strategies as a new code generation approach called DomCoder. Experimental results show that all strategies of DomCoder improve the effectiveness of domain-specific code generation under certain settings. The results also show that there is still ample room for further improvement, based on which we suggest possible future works.
@@ 1. Introduction
Despite their remarkable performance, current LLMs for code generation heavily rely on enormous training data to acquire a broad spectrum of open-domain knowledge. Notably, the evaluation of present LLMs typically revolves around open-domain benchmarks like HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021), which primarily consist of programming contests (e.g., sorting, dynamic programming), and while they showcase the capabilities of LLMs in certain aspects, they do not fully represent the intricacies and challenges associated with real-world code generation scenarios (Liu et al., 2024; Jin et al., 2024).
In this paper, we explore a more demanding code generation scenario, focusing on domain-specific code generation. We define domain-specific code as source code that is tailored specifically for and can only be applied to a particular domain (e.g., web and game), typically developed using domain-specific frameworks (e.g., HTTP, RPC, Unreal). In contrast to general-purpose code, domain-specific code poses unique challenges due to the limited availability of code corpora specific to a given domain. LLMs often struggle to acquire the necessary domain knowledge when confronted with this scarcity of domain-specific training data. This scarcity presents a significant obstacle to the task of generating domain-specific code. We aim to answer the following major questions: (1) How effective are LLMs such as ChatGPT on domain-specific code generation? (2) How to effectively prompt LLMs to produce domain-specific code? In addition, noticed that it is not always straightforward to guide LLMs on domain-specific code repositories, we also explore the following research question: (3) Can we enhance code generation models for a particular domain? Specifically, how to efficiently integrate domain knowledge into code generation models to enable them to excel in domain-specific code generation tasks?
@@ 2.3. Code Generation with Domain Libraries
Recently, code generation utilizing domain libraries has attracted research attention and there have been numerous works focused on incorporating knowledge into code generation models (Zan et al., 2022a; Liu et al., 2023b). For example, Zan et al. (Zan et al., 2022a) propose a retrieval-then-coding framework where an API retriever first identifies useful APIs and an API Coder generates code using these APIs. Shrivastava et al. (Shrivastava et al., 2023) introduce domain-specific knowledge in the prompt design process. They present the Repo-Level Prompt Generator framework, which generates example-specific prompts for LLMs. This framework incorporates context from entire repositories. Liu et al. (Liu et al., 2023b) introduce CodeGen4Libs, a technique specifically designed for library-oriented code generation. CodeGen4Libs first generates import statements and then concrete code. Zan et al. (Zan et al., 2022b) leverage the fact that library-oriented code snippets often share similar code sketches. They propose a strategy involving a sketcher and a generator trained on unlabelled data. The sketcher generates library-oriented code sketches, then the generator predicts code snippets based on them.
@@ 3.2. Evaluated Language Models
We investigate three representative LLMs: ChatGPT (AI, 2023), CodeGen (Nijkamp et al., 2022), and PolyCoder (Xu et al., 2022). They are all recently released and widely used, covering general text and code generation.
@@ 3.3. Evaluation Metrics
Therefore, in our study, we have chosen to utilize other popular metrics such as BLEU and CodeBLEU, which have also been employed by related works on code generation (Li et al., 2023b; Liu et al., 2023b). BLEU (Papineni et al., 2002) measures sequence similarity by calculating their n-gram matches. CodeBLEU (Ren et al., 2020) is a metric based on BLEU, adjusted for code evaluation, which introduces syntactic and semantic similarities into comparison. We compute both metrics using the scripts provided by the CodeXGLUE paper [^4] https://github.com/microsoft/CodeXGLUE/tree/main/Code-Code/code-to-code-trans/evaluator/CodeBLEU.
@@ 3.5.2. Results
We observe result improvement consistently on all library-specific datasets when using knowledge-enhanced prompts compared to using the plain function signature, proving that all three kinds of knowledge, including library indication, API name sequence, and API docstring can improve domain-specific code generation.
Comparing the results of basic prompts, API sequence demonstrates a significant improvement, whereas library import and docstring exhibit relatively modest enhancements. The API name sequences are closer related to the code content and have a more direct effect on code generation. In contrast, the library import prompts do carry concise and useful information, but in meager quantity. The API docstrings contain implicit knowledge in natural language, different from the target programming language, which could be indirect. Furthermore, we observed that the presence of docstrings alongside APIs does not consistently enhance performance when compared to using APIs alone, despite occasional cases where such combined prompts achieve the highest scores. This discrepancy may be attributed to the long text of the docstrings, which sometimes hinders the comprehension of APIs when incorporated into the prompt.
@@ 4.5. Results
As an essential part, we assess the accuracy of the APIs recommended by kg-GPT. This evaluation holds significance since it influences the language model’s performance in code generation relying on the suggested API sequences. We employ the same dataset outlined in Section 4.4 for training the kg-GPT. We measure the quality of API recommendations using BLEU and Hit Ratio. The Hit Ratio refers to the percentage of correctly recommended APIs out of the total number of predicted APIs. A higher Hit Ratio indicates a better performance of the API recommendation by the kg-GPT. For comparison, we provide the results of API recommendation by PolyCoder-2.7B. These results are obtained by extracting APBI sequences from the predicted code by PolyCoder. The results are presented in Table 7. We obverse that kg-GPT recommends domain API sequences with high accuracy. For example, it achieves 10 times the accuracy in the bgfx library in terms of BLEU score.

### v5

@@ Abstract.
Large language models (LLMs) such as ChatGPT have shown remarkable capabilities in code generation. Despite significant achievements, they rely on enormous training data to acquire a broad spectrum of open-domain knowledge. Besides, their evaluation revolves around open-domain benchmarks like HumanEval, which primarily consist of programming contests. Therefore, it is hard to fully characterize the intricacies and challenges associated with particular domains (e.g., web, game, and math). In this paper, we conduct an in-depth study of the LLMs in domain-specific code generation. Our results demonstrate that LLMs exhibit sub-optimal performance in generating domain-specific code, due to their limited proficiency in utilizing domain-specific libraries. We further observe that incorporating API knowledge as prompts can empower LLMs to generate more professional code. Based on these findings, we further investigate how to effectively incorporate API knowledge into the code generation process. We experiment with three strategies for incorporating domain knowledge, namely, external knowledge inquirer, chain-of-thought prompting, and chain-of-thought fine-tuning. We refer to these strategies as a new code generation approach called DomCoder. Experimental results show that all strategies of DomCoder improve the effectiveness of domain-specific code generation under certain settings.
@@ 1. Introduction
Despite their remarkable performance, current LLMs for code generation heavily rely on enormous training data to acquire a broad spectrum of open-domain knowledge. Notably, the evaluation of present LLMs typically revolves around open-domain benchmarks like HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021), which primarily consist of programming contests (e.g., sorting, dynamic programming), and while they showcase the capabilities of LLMs in certain aspects, they do not fully represent the intricacies and challenges associated with real-world code generation scenarios (Liu et al., 2024b; Jin et al., 2024).
In this paper, we explore a more demanding code generation scenario, focusing on domain-specific code generation. We define domain-specific code as source code that is tailored specifically for and can only be applied to a particular domain (e.g., web and game), typically developed using domain-specific frameworks (e.g., HTTP, RPC, Unreal). Unlike general-purpose code, domain-specific code presents distinct challenges due to the scarce availability of code corpora tailored to a specific domain. This scarcity is not just about the limited amount of training code available; rather, it highlights the relative rarity of domain-specific training data compared to the extensive pretraining data that spans multiple domains. As a result, LLMs may be short of expertise in these specific domains. This scarcity presents a significant obstacle to the task of generating domain-specific code.
We aim to answer the following major questions: (1) How effective are LLMs such as ChatGPT on domain-specific code generation? (2) How to effectively prompt LLMs to produce domain-specific code? In addition, noticed that it is not always straightforward to guide LLMs on domain-specific code repositories, we also explore the following research question: (3) Can we enhance code generation models for a particular domain? Specifically, how to efficiently integrate domain knowledge into code generation models to enable them to excel in domain-specific code generation tasks?
@@ 2.3. Code Generation with Domain Libraries
Recently, code generation utilizing domain libraries has attracted research attention and there have been numerous works focused on incorporating knowledge into code generation models (Zan et al., 2022a; Liu et al., 2023c). For example, Zan et al. (Zan et al., 2022a) propose a retrieval-then-coding framework where an API retriever first identifies useful APIs and an API Coder generates code using these APIs. Shrivastava et al. (Shrivastava et al., 2023) introduce domain-specific knowledge in the prompt design process. They present the Repo-Level Prompt Generator framework, which generates example-specific prompts for LLMs. This framework incorporates context from entire repositories. Liu et al. (Liu et al., 2023c) introduce CodeGen4Libs, a technique specifically designed for library-oriented code generation. CodeGen4Libs first generates import statements and then concrete code. Zan et al. (Zan et al., 2022b) leverage the fact that library-oriented code snippets often share similar code sketches. They propose a strategy involving a sketcher and a generator trained on unlabelled data. The sketcher generates library-oriented code sketches, then the generator predicts code snippets based on them.
@@
2.4. Studies on Code Generation with LLMs
In addition to our research, numerous empirical studies have explored code generation with LLMs (Mastropaolo et al., 2023; Liu et al., 2024a, 2023b; Vaithilingam et al., 2022). For instance, Vaithilingam et al. (Vaithilingam et al., 2022) conducted a user study to understand how programmers use and perceive LLM code generation tools. They found that while Copilot reduced the need for online searches, it introduced new challenges in code understanding, editing, and debugging. Mastropaolo et al. (Mastropaolo et al., 2023) evaluated the robustness of existing code-generation LLMs, discovering that developers using different wordings to describe the same code received varying recommendations. Liu et al. (Liu et al., 2023b) hypothesized that existing benchmarks, which rely on curated synthesis problems and test cases, may be insufficient for fully assessing the functional correctness of generated code. They proposed EvalPlus, a code synthesis evaluation framework designed to rigorously benchmark the functional correctness of LLM-generated code . More recently, Liu et al. (Liu et al., 2024a) conducted a systematic empirical assessment of the quality of code generation using ChatGPT, revealing potential issues and limitations in ChatGPT-based code generation.
Our research stands out from previous studies in both its scope and methodologies. Whereas many related studies focus on evaluating the quality and usability of generated code and challenge existing benchmarks, our work emphasizes a crucial aspect of code language models: their proficiency in generating domain-specific code. Specifically, we concentrate on how well LLMs acquire and utilize domain-specific libraries and their overall effectiveness in applying domain knowledge.
@@ 3.2. Evaluated Language Models
We investigate three LLMs: ChatGPT (AI, 2023), CodeLlama (Roziere et al., 2023), and PolyCoder (Xu et al., 2022). They are all widely used for code generation.
@@ 3.3. Evaluation Metrics
Therefore, in our study, we have chosen to utilize other popular metrics such as BLEU and CodeBLEU, which have also been employed by related works on code generation (Li et al., 2023b; Liu et al., 2023c). BLEU (Papineni et al., 2002) measures sequence similarity by calculating their n-gram matches. CodeBLEU (Ren et al., 2020) is a metric based on BLEU, adjusted for code evaluation, which introduces syntactic and semantic similarities into comparison. We compute both metrics using the scripts provided by the CodeXGLUE paper [^4] https://github.com/microsoft/CodeXGLUE/tree/main/Code-Code/code-to-code-trans/evaluator/CodeBLEU.
@@ 3.5.2. Results
We observe result improvement on library-specific datasets when using knowledge-enhanced prompts such as library indication and API name sequence compared to using the plain function signature. Comparing the results of basic prompts, API sequence demonstrates a significant improvement, whereas library import and docstring exhibit relatively modest enhancements. The API name sequences are closer related to the code content and have a more direct effect on code generation. In contrast, the library import prompts do carry concise and useful information, but in meager quantity. The API docstrings contain implicit knowledge in natural language, different from the target programming language, which could be indirect. Furthermore, we observed that the presence of docstrings alongside APIs does not consistently enhance performance when compared to using APIs alone, despite occasional cases where such combined prompts achieve the highest scores. This discrepancy may be attributed to the long text of the docstrings, which sometimes hinders the comprehension of APIs when incorporated into the prompt.
@@ 4.5. Results
As an essential part, we assess the accuracy of the APIs recommended by kg-GPT. This evaluation holds significance since it influences the language model’s performance in code generation relying on the suggested API sequences. We employ the same dataset outlined in Section 4.4 for training the kg-GPT. We measure the quality of API recommendations using BLEU and Hit Ratio. The Hit Ratio refers to the percentage of correctly recommended APIs out of the total number of predicted APIs. A higher Hit Ratio indicates a better performance of the API recommendation by the kg-GPT. For comparison, we provide the results of API recommendation by PolyCoder-2.7B. These results are obtained by extracting API sequences from the predicted code by PolyCoder. The results are presented in Table 7. We obverse that kg-GPT recommends domain API sequences with high accuracy. For example, it achieves 10 times the accuracy in the bgfx library in terms of BLEU score.